Walter is 8 years older than Ray. In 6 years, 5 times Walter’s age will equal 9 times Ray’s age. How old is Walter at present?
Answer by addingup(3677)

W = R + 8
5(W+6) = 9(R + 6)
5(R + 8 + 6) = 9R + 54
5R + 40 + 30 = 9R + 54
70 - 54 = 4R
4R = 16
R = 4
and Walter:
W = R + 8, 4 + 8 = 12
----------------------------------
Check:
W = R + 8, 4 + 8 = 12
5(W + 6) = 9(R + 6)
5(12 + 6) = 9(4 + 6)
60 + 30 = 36 + 54
90 = 90 Correct